OpenStudy (anonymous):

The formula n = rt gives the number n of boxes packed in time t at rate r. A machine packs boxes at a constant rate of 11.5 boxes per minute. How many boxes will the machine pack in 6 min?

OpenStudy (anonymous):

@Queen_Elsa321

OpenStudy (anonymous):

@uChezzy @perl

OpenStudy (phi):

do you know that n = rt means n is r * t ("r" times "t") they tell you r. what is it ? they say How many boxes will the machine pack in 6 min? do you see "t" (time) is 6 ?

OpenStudy (phi):

they tell you r. what is it ?

OpenStudy (anonymous):

11.5?

OpenStudy (phi):

yes. and t is 6 now you need to figure out r times t

OpenStudy (anonymous):

69?

OpenStudy (phi):

6*11.5 = 69 yes

OpenStudy (anonymous):

ok Thank you and fanned and medaled

OpenStudy (phi):

the idea is they give you a formula n = r t which says, to find n, multiply r times t if you can find the number for r and t, then you can do the problem. in this problem, they tell you r is 11.5 and t is 6 so you do n= 6*11.5 or n= 69